# Aa Article Title

### v1

@@ Abstract
We proposed the tensor-input tree (TT) method for scalar-on-tensor and tensor-on-tensor regression problems. We first address scalar-on-tensor problem by proposing scalar-output regression tree models whose input variable are tensors (i.e., multi-way arrays). We devised and implemented fast randomized and deterministic algorithms for efficient fitting of scalar-on-tensor trees, making TT competitive against tensor-input GP models (Sun et al., 2023; Yu et al., 2018). Based on scalar-on-tensor tree models, we extend our method to tensor-on-tensor problems using additive tree ensemble approaches. Theoretical justification and extensive experiments on real and synthetic datasets are provided to illustrate the performance of TT.
@@ 1.1 Problem and Settings
In recent years, the intersection of tensor data analysis and non-parametric modeling (Guhaniyogi et al., 2017; Papadogeorgou et al., 2021; Wang and Xu, 2024) has garnered considerable interest among mathematicians and statisticians. Non-parametric tensor models have the potential to handle complex multi-dimensional data (Bi et al., 2021) and represent spatial correlation between entries of data. This paper addresses both scalar-on-tensor (i.e., to predict a scalar response based on a tensor input) and tensor-on-tensor (i.e., both the input and output are tensors) non-linear regression problems using recursive partitioning methods, often referred to as tree(-based) models.
Supervised learning on tensor data, such as tensor regression, has significant relevance due to the proliferation of multi-dimensional data in modern applications. Tensor data naturally arises in various fields such as imaging (Wang and Xu, 2024), neuroscience (Li et al., 2018), and computer vision (Luo and Ma, 2023), where observations often take the form of multi-way arrays. Traditional regression models typically handle vector inputs and outputs, and thus can fail to capture the structural information embedded within tensor data.
@@
Existing tree regression methods such as CART, random forest, and boosting, however, do not incorporate the multi-array nature of tensor data. We aim to fill this gap by developing regression tree models tailored for tensor input and output data, enabling non-linear and potentially non-parametric modeling approaches that can capture complex interactions between elements of the tensors. First, we address scalar-on-tensor regression by proposing new scalar-output regression tree algorithms equipped with tree building strategies and loss functions tailored for tensor inputs that incorporate available low-rank tensor approximation methods. We then address the tensor-on-tensor problem by constructing an additive tree ensemble model analogous to that for tree boosting in regression (Chipman et al., 1998; Denison et al., 1998; Friedman, 2001; Friedman et al., 2004). Our tensor tree boosting uses our scalar-on-tensor tree models as “weak” learners to achieve competitive predictive performance. We show that the tree ensemble approach is particularly effective for complex outcome spaces such as in tensor-response regression.
In addition to developing methodology, we also address the computational scalability and basic theoretical guarantees of our algorithms. This is particularly important in the ensemble approach, as a large number of single trees are trained. We design fast algorithms for fitting tensor tree models that combine a divide-and-conquer strategy with approximate tree expansion algorithms. In terms of computational scalability (in big O complexity), our algorithms are competitive with existing tensor regression approaches such as tensor Gaussian processes (Yu et al., 2018; Sun et al., 2023), and alternative non-parametric models. Non-parametric tree-based models provide appropriate data partitions in conjunction with ensemble methods to extend these models into the domain of tensor inputs and outputs, integrating methodological ideas and theories from tensor decompositions and low-rank approximations in parametric tensor regression models (Zhou et al., 2013; Li and Zhang, 2017; Li et al., 2018). Our proposed non-pararmetric regression method of tensor-input trees (TT) allows us to capture multi-way dependence expressed in tensor covariates, and presents a scalable model for potentially heterogeneous tensors. In contrast to smoothing methods, TT is particularly suitable when there are non-smooth or change-of-pattern behavior in the tensor.
@@ 1.2 Regression Trees Revisited
where $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a real-valued function we wish to estimate and the $\epsilon_{i}$s are independent mean zero noises.
A single regression tree assumes a vector input in $\mathbb{R}^{d}$ and is built by recursively partitioning (Breiman et al., 1984; Hastie et al., 2009) the input space into disjoint regions and then fitting a regression mean model $m_{j}$ in each region. Specifically, suppose that the model partitions the input space into $J$ distinct and non-overlapping regions, $R_{1},R_{2},\ldots,R_{J}\subset\mathbb{R}^{d},R_{j}\cap R_{k}=\emptyset$ when $j\neq k$. The regression tree model can be written as $y_{i}=g(X_{i})+\epsilon_{i}$, where the conditional mean function $g$ in this regression model is defined as:
$g(X;\mathcal{T},\mathcal{M})=\sum_{j=1}^{J}m_{j}(X)\cdot I(X\in R_{j}),$ (2)
where $I(X\in R_{j})$ is an indicator function, $\mathcal{T}$ is the tree structure that determines the partition of the input space, and $\mathcal{M}$ determines the tree model’s prediction values. For any region $R_{j}$, the model predicts the same value over the entire region; this value is often the sample mean of the responses of the training data in the region, i.e., $m_{j}(X)=\hat{y}_{R_{j}}$ where $\hat{y}_{R_{j}}=\frac{1}{N_{j}}\sum_{i=1}^{n}y_{i}\cdot I(X_{i}\in R_{j}),N_{j}=\sum_{i=1}^{n}I(X_{i}\in R_{j})$. The input space is partitioned by minimizing a splitting criterion such as the sum of squared residuals (or sum of variances)of responses within each region:
@@
Typically $J=2$ so that each tree split induces two (sub)regions corresponding to left and right children nodes, in which case we can use the vector notation $\bm{y}=(y_{1},\cdots,y_{n}),I\subset\{1,\cdots,n\}$ and $\bm{y}[I,:]=\{y_{i}\mid i\in I\}$:
$\text{SSE}(j_{0},j_{1})=\sum_{j=1}^{2}\frac{1}{N_{j}}\sum_{I_{j}=\{k\mid\bm{X}[k,j_{1}]\in R_{j}\}}\left\|\bm{y}[I_{j},:]-\frac{1}{|I_{j}|}\bm{1}_{|I_{j}|}^{\top}\bm{y}[I_{j},:]\cdot\bm{1}_{|I_{j}|}\right\|_{2}^{2}$ (4)
where the regions $R_{1}$ and $R_{2}$ are defined by bisecting a given region along the $j_{1}$th axis at the observed input $\bm{X}[j_{0},j_{1}]$, i.e., at the $(j_{0},j_{1})$-th entry in the design matrix $\bm{X}\in\mathbb{R}^{n\times d}$ created by stacking the vectors $X_{1},\cdots,X_{n}$ as rows. We adopt a top-down greedy algorithm using criterion (3) (see Algorithm 2) for fitting a regression tree model (2) at each of its leaf nodes.
Bayesian regression trees (Chipman et al., 1998; Denison et al., 1998; Chipman et al., 2010, 2012) extend this approach by placing a prior distribution on the space of possible trees and their hyper-parameters like splitting coordinates and values, hence providing uncertainty quantification over predictions. Posterior distributions of the trees are then explored and updated using Markov chain Monte Carlo (MCMC) methods. The Bayesian approach allows for full posterior inference, including point and interval estimates of the unknown regression function and the effects of potential predictors (Pratola et al., 2014; Pratola, 2016; Luo and Pratola, 2022).
Boosting is an ensemble learning technique that enhances model performance by combining multiple weak learners, usually decision trees, into a strong learner. This method sequentially trains base models, with each new model correcting the errors of the previous ones, thus reducing bias and variance. Pioneering boosting algorithms include AdaBoost (Freund and Schapire, 1997), which adjusts weights of misclassified instances, and Gradient Boosting (Friedman, 2001; Friedman et al., 2004), which fits new models to residual errors. Bayesian Additive Regression Trees (BART), introduced by Chipman et al. (2010), combine Bayesian methods with boosting principles, modeling responses as a sum of regression trees and providing a probabilistic framework with uncertainty quantification. BART is a non-parametric Bayesian regression model which enables full posterior inference on an additive ensemble of regression trees. This paper will focus on non-Bayesian approaches of ensemble construction, but generalization to Bayesian models is natural in most scenarios (Wang and Xu, 2024).
@@ 1.3 Tensor Input Linear Regression
A key ingredient of tree-based regression models is the baseline model that characterizes the conditional response distribution within the terminal nodes of the recursive partition (i.e., the leaves of the tree). Instead of adopting a constant baseline, which ignores any further dependence between the outcome and predictors, our base model will use existing linear regression models tailored for tensor inputs. The more flexible baseline model will not only improve predictive performance, but it will also generally lead to more parsimonious trees and therefore improve computational efficiency.
@@
Next we briefly review the relevant tensor-input linear models that we consider later as the base models. When the regression function $f$ in (1) is assumed to be linear, we can fit a vector input linear regression model for estimation and prediction of the output $y$. In matrix product notation, one can express this model as $\bm{y}=\bm{X}\bm{\beta}+\bm{\epsilon}$ with $\bm{X}\in\mathbb{R}^{n\times d}$ and $\bm{y}\in\mathbb{R}^{n\times 1}$. We can estimate the model by solving the loss function (i.e., $L_{2}$ residual) $\|\bm{y}-\bm{X}\bm{\beta}\|_{2}$ for the coefficient $\bm{\beta}\in\mathbb{R}^{d\times 1}$. This can be generalized to tensor input (multi-)linear regression (i.e., scalar-on-tensor (multi-)linear regression) with a tensor input and a scalar response (Zhou et al., 2013; Luo and Ma, 2023). Without loss of generality, for notational simplicity we consider data pairs consisting of a 3-way input tensor $\bm{X}_{i}\in\mathbb{R}^{1\times d_{1}\times d_{2}}$, where the first dimension is treated as an observation index as in vector input linear models, and a scalar response $y_{i}\in\mathbb{R}^{1}$. This setup can directly handle tensor inputs via the tensor product $\circ$ in the following tensor regression model:
$y_{i}$ $=\bm{X}_{i}\circ\bm{B}+E_{i},i=1,\cdots,n\text{ or in tensor format, }$, (5)
$\bm{y}$ $=\bm{X}^{(n)}\circ\bm{B}^{(n)}+\bm{E},$ (6)
@@
where the matrix $\bm{y}\in\mathbb{R}^{n\times 1}$ stacks the responses $y_{1},\ldots,y_{n}$, the tensor $\bm{X}^{(n)}\in\mathbb{R}^{n\times d_{1}\times d_{2}}$ stacks the inputs $\bm{X}_{1},\ldots,\bm{X}_{n}$ along the first dimension, the tensor $\bm{B}\in\mathbb{R}^{d_{1}\times d_{2}}$ collects the regression coefficients, and the tensor $\bm{E}\in\mathbb{R}^{n\times 1}$ stacks the i.i.d. Gaussian noise $E_{i}\in\mathbb{R}^{1}$. The tensor products are defined as in Zhou et al. (2013). As in the vector input case, the solution to (6) corresponds to the least square problem $\min_{\bm{B}^{(n)}}\|\bm{y}^{(n)}-\bm{X}^{(n)}\circ\bm{B}^{(n)}\|_{2}$. In this paper, the superscript $(n)$ will denote an object created by stacking $n$ quantities (corresponding to $n$ observations) along the first dimension. If it is obvious that the object is stacked, we will omit the superscript to reduce visual clutter.
We use the state-of-the-art $\mathtt{tensorly}$ (Kossaifi et al., 2019) implementation for scalar-on-tensor CP/Tucker regressions for leaf models in our tree regression method.
@@
As Zhou et al. (2013) and Li et al. (2018) noted, scalar-on-tensor linear regression can extend to $D$-mode tensor inputs $\bm{X}\in\mathbb{R}^{n\times d_{1}\times d_{2}\times\cdots\times d_{D}}$ with more complex notations. We focus on $D=3$ and $D=4$ (i.e., $\bm{X}\in\mathbb{R}^{n\times d_{1}\times d_{2}}$ and $\bm{X}\in\mathbb{R}^{n\times d_{1}\times d_{2}\times d_{3}}$), which are common in applications, with notation centered on $D=3$. The model (5) has high-dimensional coefficients $\mathbf{B}$ ($n\times d_{1}\times d_{2}$), complicating model fitting and optimization. Current state-of-the-art implementations like CatBoost (Prokhorenkova et al., 2018) and XGBoost (Chen and Guestrin, 2016) support up to $D=2$ and overlook spatial correlation in $\bm{X}\circ\bm{B}$.
@@
Tensor decomposition (Kolda and Bader, 2009; Johndrow et al., 2017) is a popular approach to reducing the dimensionality and hence model complexity by aiming to capture multi-way interactions in the data using lower-dimensional representations. In particular, CP and Tucker decompositions extend the idea of low-rank approximation (e.g., SVD) from matrices to tensors, offering unique ways to explore and model the possible sparse structure in tensors. CP decomposition represents a tensor $\bm{T}\in\mathbb{R}^{t_{1}\times t_{2}\times t_{3}}$ (specifically $\bm{X}\in\mathbb{R}^{n\times d_{1}\times d_{2}}$ in (6), $t_{1}=n$, $t_{2}=d_{1}$, $t_{3}=d_{2}$) as a linear combination of $R$ rank-$1$ tensors $\mathbf{a}_{r}\times\mathbf{b}_{r}\times\mathbf{c}_{r}$:
@@
where $\times$ denotes the outer product (i.e., Kronecker product) between vectors, and $\lambda_{r}$ is the scalar weight of the $r$-th component for factor vectors $\mathbf{a}_{r}\in\mathbb{R}^{t_{1}},\mathbf{b}_{r}\in\mathbb{R}^{t_{2}},\mathbf{c}_{r}\in\mathbb{R}^{t_{3}}$, reducing the number of coefficients in $\bm{B}$ from $nd_{1}d_{2}$ to $R(1+t_{1}+t_{2}+t_{3})\leq R(1+n+d_{1}+d_{2})$. Tucker decomposition represents the tensor $\bm{T}$ (specifically $\bm{X}\in\mathbb{R}^{n\times d_{1}\times d_{2}}$ in (6), $t_{1}=n$, $t_{2}=d_{1}$, $t_{3}=d_{2}$) using a (usually dense) core tensor $\mathbf{G}\in\mathbb{R}^{R_{1}\times R_{2}\times R_{3}}$, and factor matrices $\mathbf{A}_{1}\in\mathbb{R}^{R_{1}\times t_{1}},\mathbf{A}_{2}\in\mathbb{R}^{R_{2}\times t_{2}},\mathbf{A}_{3}\in\mathbb{R}^{R_{3}\times t_{3}}$ where $R_{1}\leq t_{1},R_{2}\leq t_{2},R_{3}\leq t_{3}$:
@@
Here ×$\times_{q}$ denotes the product along mode-$q$ ($q=1,2,\cdots,D$), which is the mathematical operation of multiplying a tensor by another tensor along a specific mode by permuting the mode-$q$ in front and permuting it back. That is, we pull one dimension to the front, flatten it, transform it by a matrix, then put everything back the way it was, except that the numbers along one dimension are mixed and changed (Kolda and Bader, 2009). The core tensor $\mathbf{G}$ has rank $(R_{1},R_{2},R_{3})$ along each mode but is typically much smaller in size than $\bm{T}$. We use the second equivalent summation as in (3) of Li et al. (2018) for $\lambda_{r_{1},r_{2},r_{3}}\in\mathbb{R},\mathbf{a}_{r_{1}}\in\mathbb{R}^{t_{1}},\mathbf{b}_{r_{2}}\in\mathbb{R}^{t_{2}},\mathbf{c}_{r_{3}}\in\mathbb{R}^{t_{3}}$, reducing the number of coefficients parameters from $nd_{1}d_{2}$ to $(1+t_{1}R_{1}+t_{2}R_{2}+t_{3}R_{3})\leq(1+nR+d_{1}R+d_{2}R)$ when we use the same rank $R$ along each mode.
For tensor regression (5), we adopt a low-rank decomposition for coefficient $\bm{B}$ in (6) using CP or Tucker decompositions and use the factor matrices through the linear model (6).
Performance Metrics
1.4 Organization
The paper is organized as follows: Section 2 introduces the ingredients for fitting a single tree model with tensor inputs. Section 2.1 introduces splitting criteria for tensor inputs, followed by algorithm designs for reducing the computational complexity when splitting based on these criteria. The complexity measures of the models needed for pruning the tree structure are discussed in Section 2.3, and the grow-prune fitting procedure for our TT model is summarized in Section 2.4. Section 3 begins by reviewing ensemble techniques for improving single trees (Section 3.1), and then introduces entry-wise and low-rank methods for handling tensor output using ensembles (Section 3.2). Section 4 provides two groups of theoretical results including the consistency of leaf models and the oracle bounds for predictions. Section 5 investigates the effect of using novel splitting criteria (Section 5.1), compares to other tensor models in terms of prediction and efficiency (Section 5.2), and concludes with a tensor-on-tensor application example (Section 5.3). Section 6 provides discussions and potential future directions based on this model.

### v2

# SMSM References

@@ Abstract
We proposed the tensor-input tree (TT) method for scalar-on-tensor and tensor-on-tensor regression problems. We first address scalar-on-tensor problem by proposing scalar-output regression tree models whose input variables are tensors (i.e., multi-way arrays). We devised and implemented fast randomized and deterministic algorithms for efficient fitting of scalar-on-tensor trees, making TT competitive against tensor-input GP models (Sun et al., 2023; Yu et al., 2018). Based on scalar-on-tensor tree models, we extend our method to tensor-on-tensor problems using additive tree ensemble approaches. Theoretical justification and extensive experiments, including testing robustness to entrywise input tensor noise, are provided on real and synthetic datasets to illustrate the performance of TT.
@@
Keywords: Decision tree regressions, scalar-on-tensor regressions, tensor-on-tensor regressions, ensemble methods.
@@ 1.1 Problem and Settings
In recent years, the intersection of tensor data analysis and non-parametric modeling methods (Guhaniyogi et al., 2017; Papadogeorgou et al., 2021; Wang and Xu, 2024) has garnered considerable interest among mathematicians and statisticians. Non-parametric tensor models have the potential to handle complex multi-dimensional data (Bi et al., 2021) and represent spatial correlation between entries of data. This paper addresses both scalar-on-tensor (i.e., to predict a scalar response based on a tensor input) and tensor-on-tensor (i.e., both the input and output are tensors) non-linear regression problems using recursive partitioning methods, often referred to as tree(-based) models.
Supervised learning on tensor data, such as tensor regression, has significant relevance due to the proliferation of multi-dimensional data in modern applications. Tensor data naturally arises in various fields such as imaging (Wang and Xu, 2024), neuroscience (Li et al., 2018), and computer vision (Luo and Ma, 2025), where observations often take the form of multi-way arrays. Traditional regression models typically handle vector inputs and outputs, and thus can fail to capture the structural information embedded within tensor data.
@@
Existing tree regression methods like CART, random forest, and boosting do not handle tensor data’s multi-array characteristics. We develop regression tree models for tensor inputs and outputs to enable non-linear and possibly non-parametric modeling that captures complex tensor interactions. We introduce scalar-on-tensor regression with new scalar-output tree algorithms featuring strategies and loss functions designed for tensor data, integrating low-rank tensor approximation methods.
We then address the tensor-on-tensor problem by constructing an additive tree ensemble model analogous to that for tree boosting in regression (Chipman et al., 1998; Denison et al., 1998; Friedman, 2001; Friedman et al., 2004). Our tensor tree boosting uses our scalar-on-tensor tree models as “weak” learners to achieve competitive predictive performance. We show that the tree ensemble approach is particularly effective for complex outcome spaces such as in tensor-response regression.
In addition to developing methodology, we also address the computational scalability and basic theoretical guarantees of our algorithms. This is particularly important in the ensemble approach, as a large number of single trees are trained. Our algorithms are competitive with existing tensor regression approaches such as tensor Gaussian processes (Yu et al., 2018; Sun et al., 2023), and alternative non-parametric models. Non-parametric tree-based models provide appropriate data partitions in conjunction with ensemble methods to extend these models into the domain of tensor inputs and outputs, integrating methodological ideas and theories from tensor decompositions and low-rank approximations in parametric tensor regression models (Zhou et al., 2013; Li and Zhang, 2017; Li et al., 2018). Our proposed non-parametric regression method of tensor-input trees (TT) allows us to capture multi-way dependence expressed in tensor covariates, and presents a scalable model for potentially heterogeneous tensors. In contrast to smoothing methods, TT is particularly suitable when there are non-smooth or change-of-pattern behavior in the tensor.
In connection with generalized tree-based models with inputs in Fréchet spaces (Capitaine et al., 2024) (a.k.a., Fréchet trees in Qiu et al. (2024)), the tensor space can be viewed as a Fréchet space with additional structure that admits unique moments (Luo and Ma, 2025) and low-rank decompositions (Kolda and Bader, 2009). This allows us to introduce novel tensor-specific splitting rules (i.e., LAE and LRE in Sec 2.1), whereas existing works (Capitaine et al., 2024; Qiu et al., 2024) only consider a variance-based splitting criteria using clustering following the spirit of honest forest (Athey et al., 2019). In light of this, Krawczyk (2021) developed a clustering-based splitting tensor-input tree model. To reduce the computational burden introduced by solving Fréchet moments in splitting, Bulté and Sørensen (2024) propose to solve the optimization among a smaller fraction of solution space called Fréchet Medoid, which is very similar to our subsampling method, but different from our branch-and-bound strategy. In addition to single tensor-input trees, we also leverage boosting ensembles (Friedman, 2001) in order to reduce bias in tensor-output cases; in contrast, a random forest approach leaves the bias unchanged and instead aims to reduce variance.
@@ 1.2 Regression Trees Revisited
where $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a real-valued function and the $\epsilon_{i}$s are independent mean zero noises.
A single regression tree assumes a vector input in $\mathbb{R}^{d}$ and is built by recursively partitioning (Breiman et al., 1984; Hastie et al., 2009) the input space into disjoint regions $R_{1},R_{2},\ldots,R_{J}\subset\mathbb{R}^{d},R_{j}\cap R_{k}=\emptyset$ when $j\neq k$, before fitting a regression mean model $m_{j}$ in each region. The regression tree model can be written as
$y_{i}=g(X_{i})+\epsilon_{i},\qquad g(X;\mathcal{T},\mathcal{M})=\sum_{j=1}^{J}m_{j}(X)\cdot I(X\in R_{j}),$ (2)
where $I(X\in R_{j})$ is an indicator function, $\mathcal{T}$ is the tree structure that dictates the partition of the input space, and $\mathcal{M}$ determines the tree model’s prediction values. The partition is created by minimizing a splitting criterion such as the sum of squared residuals (or sum of variances) of responses within each region:
@@
Typically $J=2$ so that each tree split creates left and right children nodes which correspond to two (sub)regions defined by bisecting a given region along the $j_{1}$th axis at the observed $\bm{X}[j_{0},j_{1}]$, i.e., at the $(j_{0},j_{1})$-th entry in the design matrix $\bm{X}\in\mathbb{R}^{n\times d}$ created by stacking the vectors $X_{1},\cdots,X_{n}$ as rows. We adopt a top-down greedy algorithm using criterion (3) (see Algorithm 2) to create the partition. The model then assigns a prediction value to each region; this value is often the sample mean of the responses of the training data in the region, i.e., $m_{j}(X)=\hat{y}_{R_{j}}$ where $\hat{y}_{R_{j}}=\frac{1}{N_{j}}\sum_{i=1}^{n}y_{i}\cdot I(X_{i}\in R_{j})$ and $N_{j}=\sum_{i=1}^{n}I(X_{i}\in R_{j})$.
While there exists Bayesian regression tree methods (Chipman et al., 1998; Denison et al., 1998; Chipman et al., 2010, 2012), this paper will focus on non-Bayesian approaches of ensemble construction, but generalizing our model to Bayesian models is natural in most scenarios (Wang and Xu, 2024).
@@ 1.3 Tensor Input Linear Regression
Instead of adopting a constant baseline, which ignores any further dependence between the outcome and predictors, our base model will use existing linear regression models tailored for tensor inputs. The more flexible baseline model will not only improve predictive performance, but it will also generally lead to more parsimonious trees and therefore improve computational efficiency.
@@
Next we briefly review the relevant tensor-input linear models considered later as base models. When the regression function $f$ in (1) is assumed to be linear, we can fit a vector-input linear regression model for estimation and prediction of the output $y$. In matrix product notation, this model is $\bm{y}=\bm{X}\bm{\beta}+\bm{\epsilon}$ with $\bm{X}\in\mathbb{R}^{n\times d}$ and $\bm{y}\in\mathbb{R}^{n\times 1}$. We can estimate the model by solving the loss function (i.e., $L_{2}$ residual) $\|\bm{y}-\bm{X}\bm{\beta}\|_{2}$ for the coefficient $\bm{\beta}\in\mathbb{R}^{d\times 1}$. This can be generalized to tensor-input (multi-)linear regression with a tensor input and a scalar response (Zhou et al., 2013; Luo and Ma, 2025). Without loss of generality, we consider data pairs consisting of a 3-way input tensor $\bm{X}_{i}\in\mathbb{R}^{1\times d_{1}\times d_{2}}$, where the first dimension is an observation index as in vector-input models, and a scalar response $y_{i}\in\mathbb{R}^{1}$. This setup can directly handle tensor inputs via the tensor product $\circ$ in the tensor regression model:
$y_{i}=\bm{X}_{i}\circ\bm{B}+E_{i},\quad i=1,\cdots,n,\quad$ (4)
@@
where the matrix $\bm{y}\in\mathbb{R}^{n\times 1}$ stacks the responses $y_{1},\ldots,y_{n}$, the tensor $\bm{B}\in\mathbb{R}^{d_{1}\times d_{2}}$ collects the regression coefficients, and the tensor $\bm{E}\in\mathbb{R}^{n\times 1}$ stacks the i.i.d. Gaussian noise $E_{i}\in\mathbb{R}^{1}$. The tensor products are defined as $\bm{X}\circ\bm{B}=\sum_{i}^{n}\sum_{j_{1}}^{d_{1}}\sum_{j_{2}}^{d_{2}}\bm{X}[i,j_{1},j_{2}]\bm{B}[j_{1},j_{2}]$ in Zhou et al. (2013); Guo et al. (2011) (Eq. (5)) but the assumed model (4) is compatible with other tensor products as well. As in the vector-input case, the solution to (4) corresponds to the least-squares problem $\min_{\bm{B}^{(n)}}\|\bm{y}^{(n)}-\bm{X}^{(n)}\circ\bm{B}^{(n)}\|_{2}$. In this paper, the superscript $(n)$ will denote an object created by stacking $n$ quantities (corresponding to $n$ observations) along the first dimension. For $\bm{X}^{(n)}$ we stack all $n$ samples; for $\bm{B}^{(n)}$ the shared coefficient $\bm{B}$ is simply repeated $n$ times. If it is obvious that the object is stacked, we will omit the superscript.
@@
As Zhou et al. (2013) and Li et al. (2018) noted, scalar-on-tensor linear regression can extend to $D$-mode tensor inputs $\bm{X}\in\mathbb{R}^{n\times d_{1}\times d_{2}\times\cdots\times d_{D}}$ with more complex notations. We focus on $D=3$ and $D=4$ (i.e., $\bm{X}\in\mathbb{R}^{n\times d_{1}\times d_{2}}$ and $\bm{X}\in\mathbb{R}^{n\times d_{1}\times d_{2}\times d_{3}}$), which are common in applications, with notation centered on $D=3$. The model (4) has high-dimensional coefficients $\bm{B}$ complicating model fitting and optimization. Current state-of-the-art implementations like CatBoost (Prokhorenkova et al., 2018) and XGBoost (Chen and Guestrin, 2016) support up to $D=2$ and overlook spatial correlation in $\bm{X}\circ\bm{B}$. We use the state-of-the-art $\mathtt{tensorly}$ (Kossaifi et al., 2019) implementation for scalar-on-tensor CP/Tucker regressions for leaf models in our tree regression method.
@@
Tensor decomposition (Kolda and Bader, 2009; Johndrow et al., 2017) is a popular approach to reduce dimensionality and hence model complexity by aiming to capture multi-way interactions in the data using lower-dimensional representations. In particular, CP and Tucker decompositions extend the idea of low-rank approximation (e.g., SVD) from matrices to tensors, offering unique ways to explore and model the possible sparse structure in tensors. CP decomposition represents a tensor $\bm{T}\in\mathbb{R}^{t_{1}\times t_{2}\times t_{3}}$ as a linear combination of $R$ rank-$1$ tensors $\mathbf{a}_{r}\times\mathbf{b}_{r}\times\mathbf{c}_{r}$:
@@
where $\times$ denotes the outer product (i.e., Kronecker product) between vectors, and $\lambda_{r}$ is the scalar weight of the $r$-th component for factor vectors $\mathbf{a}_{r}\in\mathbb{R}^{t_{1}},\mathbf{b}_{r}\in\mathbb{R}^{t_{2}},\mathbf{c}_{r}\in\mathbb{R}^{t_{3}}$, reducing the number of coefficients in $\bm{B}$ from $nd_{1}d_{2}$ to $R(1+t_{1}+t_{2}+t_{3})\leq R(1+n+d_{1}+d_{2})$. Tucker decomposition represents the tensor $\bm{T}$ using a (usually dense) core tensor $\mathbf{G}\in\mathbb{R}^{R_{1}\times R_{2}\times R_{3}}$, and factor matrices $\mathbf{A}_{1}\in\mathbb{R}^{R_{1}\times t_{1}},\mathbf{A}_{2}\in\mathbb{R}^{R_{2}\times t_{2}},\mathbf{A}_{3}\in\mathbb{R}^{R_{3}\times t_{3}}$ where $R_{1}\leq t_{1},R_{2}\leq t_{2},R_{3}\leq t_{3}$:
@@
Here ×$\times_{q}$ denotes the product along mode-$q$ ($q=1,2,\cdots,D$), which is the multiplication of a tensor by another tensor along a specific mode by permuting the mode-$q$ in front and permuting it back. That is, we pull one dimension to the front, flatten it, transform it by a matrix, then put everything back the way it was, except that the numbers along one dimension are mixed and changed (Kolda and Bader, 2009). The core tensor $\mathbf{G}$ has rank $(R_{1},R_{2},R_{3})$ along each mode but is typically much smaller in size than $\bm{T}$. We use the second equivalent summation as in (3) of Li et al. (2018) for $\lambda_{r_{1},r_{2},r_{3}}\in\mathbb{R},\mathbf{a}_{r_{1}}\in\mathbb{R}^{t_{1}},\mathbf{b}_{r_{2}}\in\mathbb{R}^{t_{2}},\mathbf{c}_{r_{3}}\in\mathbb{R}^{t_{3}}$, reducing the number of coefficients parameters from $nd_{1}d_{2}$ to $(1+t_{1}R_{1}+t_{2}R_{2}+t_{3}R_{3})\leq(1+nR+d_{1}R+d_{2}R)$ when we use the same rank $R$ along each mode.
For tensor regression (4), we adopt a low-rank decomposition for the coefficient $\bm{B}$ using CP or Tucker decompositions and use the factor matrices through the linear model (4), following the state-of-the-art practice in Li et al. (2018); Zhou et al. (2013).
Organization.
The rest of the paper is organized as follows: Section 2 introduces the ingredients for fitting a single tree model with tensor inputs, including a discussion of splitting criteria and complexity-based pruning. Section 3 reviews ensemble techniques for improving single trees before introducing entry-wise and low-rank methods for handling tensor output using ensembles (Section 3.1). Section 4 provides two groups of theoretical results including the consistency of leaf models and the oracle bounds for predictions. Section 5 investigates the effect of using novel splitting criteria (Section 5.1), compares to other tensor models in terms of prediction and efficiency (Section 5.2), and concludes with a tensor-on-tensor application example (Section 5.3). Section 6 provides discussions and future works. The Supplementary Material contains additional simulation experiments, including for robustness to entrywise input noise, and for comparing to existing random-forest methods.